OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. d = 13

OpenStudy (anonymous):

A=πr²

OpenStudy (bahrom7893):

r = d/2 = 13/2 A = pi * r^2

OpenStudy (anonymous):

d or diameter is 2 times the radius, so r is half of the diameter.

OpenStudy (anonymous):

In your case, you are supposed to use 3.14 for the value of pi, so you should get: A=3.14*6.5²

OpenStudy (bahrom7893):

mathdrills i think we should come back later, in chat she said she was going to go cook

OpenStudy (bahrom7893):

breakfast.. Man im hungry ima go eat too haha